OpenStudy (anonymous):

A group of college students are volunteering for Help the Homeless during their spring break. They are putting the finishing touches on a house they built. Working alone, Julian can paint a certain room in 3 hours. Mike can paint the same room in 7 hours. Write an equation that can be used to find how long it will take them working together to paint the room. Include equation

OpenStudy (anonymous):

Consider their speeds: J=1room/3hours M=1room/7hours Now, we want to combine their two speeds to paint one room: Generally, number of rooms = (J's speed)(time) + (M's speed)(time). In this case, 1 room = (J)(time) + (M)(time) 1 room = (1/3)T + (1/7) T Pull out the T: 1 = T (1/3 + 1/7) 1 = T (10/21) So T = 21/10, or T = 21/10 hours. (2 and 1/10 hours). Check: In 21/10 hours, Julian painting 1 room/3 hours finishes 7/10 of the room. In 21/10 hours, Mike painting 1 room/7 hours finishes 3/10 of the room. So in 21/10 hours, they together paint the full 10/10 of the room!